OpenStudy (anonymous):

9dc/(d^2-c^2 )-(d-c)/(d+c)

OpenStudy (anonymous):

what do you want me to do simplify?

OpenStudy (anonymous):

Subtract and simpiliy by removing a factor of 1 when possible

OpenStudy (anonymous):

idk thats basically what mine is but without the 1

OpenStudy (anonymous):

can't simplify or subtract until you get a GCD

OpenStudy (anonymous):

your GCD is going to be (d+c)(d-c)

OpenStudy (anonymous):

outkast come look at mine

OpenStudy (anonymous):

let outkast help me first

OpenStudy (anonymous):

the girl that is looking at mine isnt doing anything

OpenStudy (anonymous):

hey thts me

OpenStudy (anonymous):

oohhh I did not know

OpenStudy (anonymous):

I got the bottom

OpenStudy (anonymous):

hahd are u still around

OpenStudy (anonymous):

its a little hard to tell what the problem is, is this it? \[\frac{9dc}{d^2-c^2} - \frac{d-c}{d+c}\]

OpenStudy (anonymous):

that is it

OpenStudy (anonymous):

and whats the name of the game? simplify? honestly there isnt too much that can be done. We can combine the fractions by getting common denominators, but it will still look messy.

OpenStudy (anonymous):

subtract. simplify by removing a factor of 1 if possible

OpenStudy (anonymous):

well, noting that: \[d^2-c^2 = (d+c)(d-c)\] we can multiply the second fraction by a clever form of one, namely: \[\frac{d-c}{d-c}\] if we do that the second fraction becomes: \[\frac{d-c}{d+c}*\frac{d-c}{d-c} = \frac{d^2-2dc+c^2}{d^2-c^2}\] Now you are good to subtract the fractions since their denominators are equal.

OpenStudy (anonymous):

im here

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

what joe is here

OpenStudy (anonymous):

no need for me

OpenStudy (anonymous):

whats up?

OpenStudy (anonymous):

hey man

OpenStudy (anonymous):

yeah I like your answers better hahd

OpenStudy (anonymous):

yo hahd :)

OpenStudy (anonymous):

nm u?

OpenStudy (anonymous):

same here, its been a pretty nice chill day.

OpenStudy (anonymous):

I do not know the answer

OpenStudy (anonymous):

ill solve it dw

OpenStudy (anonymous):

yo joe when u goin to uni

OpenStudy (anonymous):

i converted the second fraction to the same denominator, its just a matter of subtraction now. @hahd i start school next wed, i cant wait lol

OpenStudy (anonymous):

-(d^2-11dc+c^2)/((d-c)(d+c)) there thts the answer

OpenStudy (anonymous):

yo joe is uni of texas residence nice?

OpenStudy (anonymous):

i say thats good, i dont really see anything else you can do with the problem =/ @hahd i wouldnt know =/ i live off campus in an apartment. From what ive seen though it looks nice.

OpenStudy (anonymous):

ah smart person all uni plan on $20000 from me if i stay on res, its ridiculous

OpenStudy (anonymous):

wow that seems ridiculous o.O

OpenStudy (anonymous):

yh but there's a prog called osap that gives u aloan u can pay it off later

OpenStudy (anonymous):

9-2 = 11 or am i looking at two different problems

OpenStudy (anonymous):

its 9-(-2)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

there is so much work bunched in so many different areas i'm like trying to pick out

OpenStudy (anonymous):

ahh indeed

OpenStudy (anonymous):

joe gives nice explanation

OpenStudy (anonymous):

ha

OpenStudy (anonymous):

HAHD ANSWER MY QUESTION NAO

OpenStudy (anonymous):

guys can u help me level up one more i wanna sleep a hero

OpenStudy (anonymous):

cant havent taken calculus yet

OpenStudy (anonymous):

ill give you an easy one

OpenStudy (anonymous):

i'll post it in five seconds and you just search p series test

OpenStudy (anonymous):

on google

OpenStudy (anonymous):

calculus isnt that great imo. there are other harder topics that are more interesting. thats just a personal opinion though.

OpenStudy (anonymous):

Likeeee?

OpenStudy (anonymous):

i personally like proofing

OpenStudy (anonymous):

Number Theory, or Linear Algebra and Matrix Theory, or Modern Abstract Algebra.

OpenStudy (anonymous):

i find it confusing at times because of hte slight changes with like variables

OpenStudy (anonymous):

i enjoy those more than Calculus. again thats just a personal opinion.

OpenStudy (anonymous):

Linear algebra.... I was planning on taing that and calc 3 next semeseter

OpenStudy (anonymous):

thanks for telling me tha it's hard

OpenStudy (anonymous):

YES THEY ARE AWESOME TOPICs, especially when u study computer science

OpenStudy (anonymous):

if its a first class in Linear Algebra its not that hard! sry i dont want to make it seems like impossible.

OpenStudy (anonymous):

@hahd yeah those topics have a wide range of applications in computer science.

OpenStudy (anonymous):

well i can switch it if i do it soon... would you say linear or differentials was harder?

OpenStudy (anonymous):

a family member of mine said that linear algebra was actually simple for her, like the matrices and stuff. But she said that her differential teacher sucked so thats why she hated it

OpenStudy (anonymous):

that depends on how the professor teaches the course. I think Linear Algebra could be harder if taught a certain way, but most likely it wont be bad.

OpenStudy (anonymous):

like, when i took Linear Algebra (first course), it wasnt bad, but then i sat through a different professors Linear Algebra course to get a better understanding, it was WAY WAY harder than my class.

OpenStudy (anonymous):

the only tough thing i find in calc 2 (which i'm in right now) is the changing of variables for stuff like H(y) when the past integrations they used another way of saying it achapter before

OpenStudy (anonymous):

haha well i will have the same teacher for both my calc 3 and linear algebra

OpenStudy (anonymous):

I think he also teaches the differential this semester too

OpenStudy (anonymous):

joe help me with the medal situation

OpenStudy (anonymous):

pls

OpenStudy (anonymous):

i wana be a hero

OpenStudy (anonymous):

i dont think your linear algebra class will be that bad then. Calc 3 is one of the more computational math classes than a theoretical math class. Linear Algebra can be taught either way, as a computational or theoretical. If your calc 3 professor is teaching Linear Alg., it will probably be computational.

OpenStudy (anonymous):

where have you posted Hahd? i'll give you medals.

OpenStudy (anonymous):

just click on my name some ppl dont even bothering givin medals after i help them